OpenStudy (anonymous):

determine theoretical probablility the product of rolling 2 dice will be greater tahn 10

OpenStudy (anonymous):

than*

OpenStudy (anonymous):

How many possible sums are there?

OpenStudy (anonymous):

1. How many possible combinations. Answer is 6*6 = 36. 2. How many of those combinations add up to 11 or 12? Answer 6+6=12 5+6=11, 6+5=11 so 3 You have a 3/36=1/12 chance of rolling a sum greater than 10 I think.